# The deep multi-FBSDE method: a robust deep learning method for coupled FBSDEs

### v2

@@ Abstract
backward stochastic differential equations (FBSDEs), focusing on cases where the deep BSDE method of Han, Jentzen, and E (2018) fails to converge. To overcome the convergence issues, we consider a family of FBSDEs that are equivalent to the original problem in the sense that they satisfy the same associated partial differential equation (PDE). Our algorithm proceeds in two phases: first, we approximate the initial condition for the FBSDE family, and second, we approximate the original FBSDE using the initial condition approximated in the first phase. Numerical experiments show that our method converges even when the standard deep BSDE method does not.
@@ 1 Introduction
Solving forward-backward stochastic differential equations (FBSDEs) using deep learning has emerged as a powerful technique to overcome the curse of dimensionality inherent in classical numerical methods. Han, Jentzen, and E’s seminal work [17] introduced the deep BSDE method, employing neural networks to approximate solutions of high-dimensional parabolic partial differential equations (PDEs) by reformulating them as FBSDEs. This breakthrough demonstrated remarkable performance on problems previously considered intractable, showcasing the ability of neural networks to efficiently handle complex nonlinearities at scales unattainable by traditional approaches. Following this success, several deep learning-based strategies have emerged, notably [6, 7, 5, 14, 28, 23, 24, 31, 19, 3], with convergence analyses provided in, e.g., [18, 21, 9, 12, 16, 22, 29, 25, 2]. Concurrently, a separate branch known as backward-type methods, closer in spirit to classical dynamic programming algorithms—has developed, see e.g., [20, 11, 13, 27, 26, 20, 15]. For a comprehensive overview of machine learning algorithms for PDE approximation, we refer to [8].
@@
Despite these advancements, forward-type deep BSDE methods, including the original method by Han et al., still face significant convergence challenges when addressing certain classes of FBSDEs, such as those encountered in stochastic control problems [1]. Specifically, these forward-type approaches often fail to converge or become trapped in suboptimal solutions due to ill-conditioned optimization landscapes [20, 1, 10]. Although backward-type methods may not suffer from precisely the same convergence issues, they encounter different challenges, especially when dealing with coupled FBSDEs. The scope of this paper focuses specifically on the improvement of forward-type methods.
@@
To address the convergence challenges, we propose a robust modification of the deep BSDE method that we name the deep multi-FBSDE method, applicable for a broad class of FBSDEs and their related PDEs. Our approach alleviates optimization difficulties by utilizing a family of FBSDEs that share the same initial condition as well as the same functional forms as the original FBSDE. In other words, all FBSDEs in this family are solved by the same PDE. We propose a two-phase optimization procedure. In the first phase, we approximate the joint initial conditions for all FBSDEs. In the second phase, we apply a reduced version of the deep BSDE method, where we optimize only over the control processes while using the initial condition obtained from the first phase. This leads to a numerical scheme which we experimentally show has consistent and accurate convergence, even for strongly coupled, nonlinear FBSDEs. Our numerical experiments demonstrate stable and accurate solutions for problems for which the classical deep BSDE methods fail.
@@
  • A robust deep FBSDE method that effectively addresses convergence issues faced by existing deep learning-based methods for strongly coupled FBSDEs.
  • A general purpose training strategy applicable to general FBSDE systems, without restrictive assumptions such as the existence of an equivalent stochastic control problem, as in [1].
@@ 2 Forward backward stochastic differential equations
3 The deep BSDE and deep multi-FBSDE methods
In this section, we first introduce the classical deep BSDE method and briefly discuss its convergence problems for coupled FBSDEs. Most importantly, we introduce our novel modification of the scheme. We use a Linear Quadratic (LQ) problem to demonstrate both the convergence problems of the deep BSDE method and the convergence of our method. The details of the LQ-problem can be found in Section 4.1.2 together complementary experimental results.
3.1 The deep BSDE method
The deep BSDE method relies on a reformulation of the FBSDE (1) into two forward SDEs, one with an a priori unknown initial value. It relies moreover on the Markov property of the FBSDE, which guarantees that $Z_{t}=\zeta^{*}(t,X_{t})$, for some function $\zeta^{*}\colon[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{k}$, that we refer to as a Markov map. Optimization is done with respect to such functions and the initial values $y_{0}$ with the objective to satisfy the terminal condition. More precisely, the FBSDE (1) is reformulated into the following variational problem
@@ 3.1 The deep BSDE method
We obtain a semi-discretized version of (3) by approximating the Itô integrals using the Euler–Maruyama scheme. We assume an equidistant temporal grid with $N\in\mathbb{N}$ time steps given by $\pi=\{t_{0},t_{1},\ldots,t_{N}\}$ with $t_{0}=0$, $t_{N}=T$ and step size $h=t_{n+1}-t_{n}$, and denote the i.i.d. Wiener increments by $\Delta W_{n}=W_{t_{n+1}}-W_{t_{n}}\sim\mathcal{N}(0,h)$. The semi-discrete version of (3) is given by:
@@
In the deep BSDE method, we search for $y_{0}\in\mathbb{R}$, a simple trainable parameter, and for $\zeta=(\zeta_{n})_{n=0}^{N-1}$ in a parametric space defined by the architecture of a specified neural network.
@@
Under conditions that guarantee the same or similar a posteriori bound, such as those in [18], the function $\mathbb{E}|\hat{Y}_{N}^{h,y_{0},\zeta}-g(X_{N}^{h,y_{0},\zeta})|^{2}$ is a highly suitable loss function and guarantees an approximation with an error bounded by the loss after training and the time step. When the assumptions are not met, the situation is less fortunate. Counterexamples of FBSDEs from stochastic control are presented in [1] for which it is possible to make the right-hand side of (5) arbitrarily small, but with the left-hand side remaining large and discrete solutions converging to severely wrong processes. In this paper, we provide four more counterexamples. Thus, while the a posteriori bound in [18] is a powerful tool under its specific conditions, extrapolating it beyond these conditions is a shaky practice.
For instructive purposes and since we address the convergence problem documented in [1], with a novel method, we briefly repeat the problem with a numerical example. In order to demonstrate that the optimization landscape for the deep BSDE method is problematic for strongly coupled FBSDEs, we use the mean-squared error (MSE)
@@
This is the MSE with $y_{0}$ fixed and only $\zeta$ being optimized. If joint optimization of $y_{0}$ and $\zeta$, as in the deep BSDE method, is supposed to yield a good approximation of the FBSDE, then by necessity the minimum of $y_{0}\mapsto\mathrm{MSE}(y_{0})$ must be close to the true $Y_{0}$ of the FBSDE. For an LQ problem, detailed in Section 4.1.2 below, Figure 1 shows that the optimization landscape $y_{0}\mapsto\mathrm{MSE}(y_{0})$ instead has a minimum $y_{0}^{*}$ very far from the true $Y_{0}$. Figure 2 shows the exact $Y$ and the output of the deep FBSDE method, pathwise and in mean. We note that although the initial condition is significantly incorrect, the terminal condition is well approximated, i.e., $g(X_{N}^{h,y_{0},\zeta})-Y_{N}^{h,y_{0},\zeta}$ is small both on the pathwise level and in the mean-squared sense. It should be noted that the numerical algorithm has converged, as the loss function is no longer decreasing and $y_{0}$ remains constant (not shown here but Figure 6 qualitatively shows the same behavior for a different problem). Moreover, the algorithm is consistent, with repeated runs yielding similar values for the loss function and $y_{0}$. The reader might wonder if a smaller time step $h$ in the approximation would improve approximation of $Y_{0}$. Figure 5 indicates a negative answer to this valid thought by showing non-convergence of our implementation of the deep BSDE method (after thorough hyperparameter optimization) (the case $K=1$ in that figure).
@@
Figure 2: The almost analytical $Y$-process and the approximation obtained by using the deep BSDE method. Left: One representative path. Right: A sample mean of size $2^{12}$.
The behavior depicted in Figures 12 is not confined to a narrow class of specific problems or special cases. For example, it is not limited to strongly coupled FBSDEs, and one can easily construct low-dimensional (e.g., $d=2$) examples of decoupled BSDEs with smooth solutions to the associated PDE, wherein the deep BSDE method fails. In particular, quadratic BSDEs appear susceptible to this phenomenon. A more detailed, although mostly qualitative and empirical, discussion of the reasons for this lack of convergence is provided in [1].
@@ 3.2 The deep multi-FBSDE method
The deep multi-FBSDE is divided into two phases. Phase I is based on approximating (9) for suitable $K$ and $\psi_{1},\dots,\psi_{K}$, with time discretization, neural network approximation of $\zeta$ and stochastic gradient descent. In the second phase, a better approximation of $\zeta$ is obtained by solving the classical deep BSDE method with the exception that only $\zeta$ is optimized and $y_{0}$ is taken from Phase I. For easy and low-dimensional problems, Phase I gives an accurate approximation of $\zeta$ and Phase II is not needed. For more challenging and high-dimensional problems, we have observed that Phase II improves the approximation accuracy. We believe that this has to do with the fact that in Phase I the Markov map $\zeta$ is optimized for good fit along typical trajectories of $K$ different SDEs, i.e., in a larger domain than in Phase II, where it is only trained along typical trajectories of the original forward SDE, i.e., for a more limited task.
We next state the scheme and similar to the deep BSDE method let $\zeta$ represent $\zeta_{0},\ldots,\zeta_{N-1}$.
@@
We demonstrate the new numerical scheme on the LQG-problem, detailed in Section 4.1.2, for which the deep BSDE method was shown to fail in Section 3.1. In Phase I, we take $K=1,2,3,4$, with the choices $\psi_{1}=0$, $(\psi_{1},\psi_{2})=(0,b_{1})$, $(\psi_{1},\psi_{2},\psi_{3})=(0,b_{1},-b_{1})$ and $(\psi_{1},\psi_{2},\psi_{3},\psi_{4})=(0,b_{1},-b_{1},-0.5b_{1})$, where the drift has the form $b(x,y,z)=b_{1}(x)+b_{2}(z)$, and plot the MSE curves in Figure 3. For $K=2,3,4$ the optimization landscape is much more satisfactory than for $K=1$, which is the deep BSDE method. In fact the minima are close to the real $Y_{0}=8.94$. To further indicate that the method is robust to the choice of $K\geq 3$ (and possibly $K=2$) and $(\psi_{1},\dots,\psi_{K})$ we take $K=3$ and rather arbitrary choices $\psi_{1}=b_{1}\circ\xi$, $\psi_{2}=-b_{1}\circ\xi$, $\psi_{3}=0$, for $\xi\in\{x,-\sin(x),x\cos(x),1-\exp(\min(1,\max(-1,x)))\}$, where $\xi$ acts coordinate-wise on vectors. We obtain the four loss surfaces in Figure 4, which, based on visual inspection, have the same minima. This robustness is a very important property because when the method is applied to a problem without a reference solution, one must be able to trust that the choice of $\psi_{1},\psi_{2},\psi_{3}$ does not need to be tuned based on comparison with a reference solution. Our results indicate a robustness in the choice of $K\geq 2$ and $(\psi_{1},\dots,\psi_{K})$, but it needs to be investigated further, particularly theoretically. The deviation between $Y_{0}$ and the optimized $y_{0}^{*}$ is the discretization error due to the time stepping. This is seen empirically in the convergence plot in Figure 5 for $\xi(x)=x$, showing the experimental convergence order one, i.e., $|Y_{0}-y_{0}^{*}(h)|=\mathcal{O}(h)$. The convergence of $(X,Y,Z)$ as processes after Phase II is shown in Section 4.1.2.
Figure 3: Four MSE curves plotted for $K=1,2,3,4$, with $\psi_{1}=0$, $(\psi_{1},\psi_{2})=(0,b_{1})$, $(\psi_{1},\psi_{2},\psi_{3})=(0,b_{1},-b_{1})$ and $(\psi_{1},\psi_{2},\psi_{3},\psi_{4})=(0,b_{1},-b_{1},-0.5b_{1})$, where $b(x,y,z)=b_{1}(x)+b_{2}(z)$. In Phase II $y_{0}^{*}$ for $K=1$ and $K=2$ are the initial conditions approximated with Phase I. Note that for $K=1$, $y_{0}^{*}=28.44$ as can be visualized in Figure 1. For visualization purposes, we have scaled the curves. For $K=1,2,3,4$, the training times are approximately 400s, 500s, 600s, and 700s, respectively.
Figure 4: Four MSE curves plotted for $\psi_{1}=b_{1}\circ\xi$, $\psi_{2}=-b_{1}\circ\xi$, $\psi_{3}=0$, for $\xi\in\{x,-\sin(x),x\cos(x),1-\exp((x\wedge 1)\vee(-1))\}$, where $\xi$ acts coordinate wise on vectors. The order coincides with that of the set.
Figure 5: Empirical convergence plot for approximations of $Y_{0}$ using the deep multi-FBSDE method Phase I. Here the case $K=1$ coincides with the original deep BSDE method.

### v3

@@ Abstract
backward stochastic differential equations (FBSDEs), focusing on cases where the deep BSDE method of Han, Jentzen, and E (2018) fails to converge. To overcome the convergence issues, we consider a family of FBSDEs that are equivalent to the original problem in the sense that they satisfy the same associated partial differential equation (PDE) and initial value. Our algorithm proceeds in two phases: first, we approximate the initial condition jointly for a small number of FBSDEs from the FBSDE family, and second, we approximate the original FBSDE using the initial condition approximated in the first phase. Numerical experiments show that our method converges even when the standard deep FBSDE method does not.
@@ 1 Introduction
Solving forward-backward stochastic differential equations (FBSDEs) using deep learning has emerged as a powerful technique to overcome the curse of dimensionality inherent in classical numerical methods. Han, Jentzen, and E’s seminal work [20] introduced the deep BSDE method, employing neural networks to approximate solutions of high-dimensional parabolic partial differential equations (PDEs) by reformulating them as FBSDEs. This breakthrough demonstrated remarkable performance on problems previously considered intractable, showcasing the ability of neural networks to efficiently handle complex nonlinearities at scales unattainable by traditional approaches. Following this success, several deep learning-based strategies have emerged, notably [6, 7, 5, 17, 31, 26, 27, 34, 22, 3], with convergence analyses provided in, e.g., [21, 24, 9, 15, 19, 25, 32, 28, 2]. Concurrently, a separate branch known as backward-type methods, closer in spirit to classical dynamic programming algorithms, has developed, see e.g., [23, 14, 16, 30, 29, 23, 18]. For a comprehensive overview of machine learning algorithms for PDE approximation, we refer to [8].
@@
Despite these advancements, forward-type deep BSDE methods, including the original method in [20], still face significant convergence challenges when addressing certain classes of FBSDEs, such as those encountered in stochastic control problems [1] and beyond. Specifically, these forward-type approaches often fail to converge or become trapped in suboptimal solutions due to ill-conditioned optimization landscapes [23, 1, 10]. Although backward-type methods may not suffer from precisely the same convergence issues, they encounter different challenges, especially implementation challenges when dealing with coupled FBSDEs (see [12, 11, 13] for further discussions of implementation challenges for backward type methods when dealing with coupled FBSDEs) and accumulated errors in general (see e.g., [30]). For FBSDEs related to stochastic control problems through dynamic programming, a robust deep FBSDE method was proposed in our paper [1]. The scope of the present paper is the introduction of another robust modification of the deep FBSDE method, the deep multi-FBSDE method. It is applicable to a broader class of FBSDEs and their PDEs, not strictly defined due to the experimental nature of our paper, but not restricted to only control problems. By robustness, we mean that the method can be optimized for the equation of interest, and not the other way around, which is sometimes seen in the deep learning-based scientific computing literature. Of course, this is an inprecise definition and no method allows convergence for any equation. Still, our method has been working on the vast majority of equations we have tested and we have deliberately chosen challenging equations. We refer to the original deep BSDE method, applied directly to coupled FBSDE without modification, as the deep FBSDE method [1, 26].
@@
Our proposed approach alleviates optimization difficulties by utilizing a family of FBSDEs that share the same initial condition as well as the same functional forms as the original FBSDE. In other words, all FBSDEs in this family are solved by the same PDE. We propose a two-phase optimization procedure. In the first phase, we approximate the joint initial conditions for all FBSDEs. In the second phase, we apply a reduced version of the deep FBSDE method, where we optimize only over the control processes while using the initial condition obtained from the first phase. This leads to a numerical scheme which we experimentally show has consistent and accurate convergence, even for strongly coupled, nonlinear FBSDEs. Our numerical experiments demonstrate stable and accurate solutions for problems for which the classical deep FBSDE methods fail.
@@
  • A robust deep FBSDE method that we empirically demonstrate solves the convergence issues faced by existing deep learning-based methods for strongly coupled FBSDEs.
  • A general purpose training strategy that is demonstrated numerically to be applicable to general FBSDE systems, without restrictive assumptions such as the existence of an equivalent stochastic control problem, as in [1].
@@ 2 Forward backward stochastic differential equations
3 The deep FBSDE and deep multi-FBSDE methods
In this section, we first introduce the classical deep FBSDE method and briefly discuss its convergence problems for coupled FBSDEs. Most importantly, we introduce our novel modification of the scheme. We use a Linear Quadratic (LQ) problem to demonstrate both the convergence problems of the deep FBSDE method and the convergence of our method. The details of the LQ-problem can be found in Section 4.1.2 together complementary experimental results.
3.1 The deep FBSDE method
The deep FBSDE method relies on a reformulation of the FBSDE (1) into two forward SDEs, one with an a priori unknown initial value. It relies moreover on the Markov property of the FBSDE, which guarantees that $Z_{t}=\zeta^{*}(t,X_{t})$, for some function $\zeta^{*}\colon[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{k}$, that we refer to as a Markov map. Optimization is done with respect to such functions and the initial values $y_{0}$ with the objective to satisfy the terminal condition. More precisely, the FBSDE (1) is reformulated into the following variational problem
@@ 3.1 The deep FBSDE method
We obtain a semi-discretized version of (3) by approximating the Itô integrals using the Euler–Maruyama scheme. We assume an equidistant temporal grid with $N\in\mathbb{N}$ time steps given by $t_{0}<t_{1}<\ldots<t_{N}$ with $t_{0}=0$, $t_{N}=T$ and step size $h=t_{n+1}-t_{n}$, and denote the i.i.d. Wiener increments by $\Delta W_{n}=W_{t_{n+1}}-W_{t_{n}}\sim\mathcal{N}(0,h)$. The semi-discrete version of (3) is given by:
@@
In the deep FBSDE method, we search for $y_{0}\in\mathbb{R}$, a simple trainable parameter, and for $\zeta=(\zeta_{n})_{n=0}^{N-1}$ in a parametric space defined by the architecture of a specified neural network.
@@
Under conditions that guarantee the same or similar a posteriori bound, such as those in [21], the function $\mathbb{E}|\hat{Y}_{N}^{h,y_{0},\zeta}-g(X_{N}^{h,y_{0},\zeta})|^{2}$ is a highly suitable loss function and guarantees an approximation with an error bounded by the loss after training and the time step. When the assumptions are not met, the situation is less fortunate. Counterexamples of FBSDEs from stochastic control are presented in [1] for which it is possible to make the right-hand side of (5) arbitrarily small, but with the left-hand side remaining large and discrete solutions converging to severely wrong processes. In this paper, we provide four more counterexamples. Thus, while the a posteriori bound in [21] is a powerful tool under its specific conditions, extrapolating it beyond these conditions is a questionable practice.
For instructive purposes and since we address the convergence problem documented in [1], with a novel method, we briefly repeat the problem with a numerical example. In order to demonstrate that the optimization landscape for the deep FBSDE method is problematic for strongly coupled FBSDEs, we use the mean-squared error (MSE)
@@
This is the MSE with $y_{0}$ fixed and only $\zeta$ being optimized. If joint optimization of $y_{0}$ and $\zeta$, as in the deep FBSDE method, is supposed to yield a good approximation of the FBSDE, then by necessity the minimum of $y_{0}\mapsto\mathrm{MSE}(y_{0})$ must be close to the true $Y_{0}$ of the FBSDE. For an LQ problem, detailed in Section 4.1.2 below, Figure 1 shows that the optimization landscape $y_{0}\mapsto\mathrm{MSE}(y_{0})$ instead has a minimum $y_{0}^{*}$ very far from the true $Y_{0}$. Figure 2 shows the exact $Y$ and the output of the deep FBSDE method, pathwise and in mean. We note that although the initial condition is significantly incorrect, the terminal condition is well approximated, i.e., $g(X_{N}^{h,y_{0},\zeta})-Y_{N}^{h,y_{0},\zeta}$ is small both on the pathwise level and in the mean-squared sense. It should be noted that the numerical algorithm has converged, as the loss function is no longer decreasing and $y_{0}$ remains constant (not shown here but Figure 6 qualitatively shows the same behavior for a different problem). Moreover, the algorithm is consistent, with repeated runs yielding similar values for the loss function and $y_{0}$. The reader might wonder if a smaller time step $h$ in the approximation would improve approximation of $Y_{0}$. Figure 5 indicates a negative answer to this valid thought by showing non-convergence of our implementation of the deep FBSDE method (after thorough hyperparameter optimization) (the case $K=1$ in that figure).
@@
Figure 2: The deep FBSDE method failing to approximate the true reference solution. Left: One representative path. Right: A sample mean of size $2^{12}$.
The behavior depicted in Figures 12 is not confined to a narrow class of specific problems or special cases. For example, it is not limited to strongly coupled FBSDEs, and one can easily construct low-dimensional (e.g., $d=2$) examples of decoupled BSDEs with smooth solutions to the associated PDE, wherein the deep FBSDE method fails. In particular, quadratic BSDEs appear susceptible to this phenomenon. A more detailed, although mostly qualitative and empirical, discussion of the reasons for this lack of convergence is provided in [1].
@@ 3.2 The deep multi-FBSDE method
The deep multi-FBSDE is divided into two phases. Phase I is based on approximating (9) for suitable $K$ and $\psi_{1},\dots,\psi_{K}$, with time discretization, neural network approximation of $\zeta$ and stochastic gradient descent. In the second phase, a better approximation of $\zeta$ is obtained by solving the classical deep FBSDE method with the exception that only $\zeta$ is optimized and $y_{0}$ is taken from Phase I. For easy and low-dimensional problems, Phase I gives an accurate approximation of $\zeta$ and Phase II is not needed. For more challenging and high-dimensional problems, we have observed that Phase II improves the approximation accuracy. We believe that this has to do with the fact that in Phase I the Markov map $\zeta$ is optimized for good fit along typical trajectories of $K$ different SDEs, i.e., in a larger domain than in Phase II, where it is only trained along typical trajectories of the original forward SDE, i.e., for a more limited task.
We next state the scheme and similar to the deep FBSDE method let $\zeta$ represent $\zeta_{0},\ldots,\zeta_{N-1}$.
@@
We demonstrate the new numerical scheme on the LQG-problem, detailed in Section 4.1.2, for which the deep FBSDE method was shown to fail in Section 3.1. In Phase I, we take $K=1,2,3,4$, with the choices $\psi_{1}=0$, $(\psi_{1},\psi_{2})=(0,b_{1})$, $(\psi_{1},\psi_{2},\psi_{3})=(0,b_{1},-b_{1})$ and $(\psi_{1},\psi_{2},\psi_{3},\psi_{4})=(0,b_{1},-b_{1},-0.5b_{1})$, where the drift has the form $b(x,y,z)=b_{1}(x)+b_{2}(z)$, and plot the MSE curves in Figure 3. For $K=2,3,4$ the optimization landscape is much more satisfactory than for $K=1$, which is the deep FBSDE method. In fact the minima are close to the real $Y_{0}=8.94$. To further indicate that the method is robust to the choice of $K\geq 3$ (and possibly $K=2$) and $(\psi_{1},\dots,\psi_{K})$ we take $K=3$ and rather arbitrary choices $\psi_{1}=b_{1}\circ\xi$, $\psi_{2}=-b_{1}\circ\xi$, $\psi_{3}=0$, for $\xi\in\{x,-\sin(x),x\cos(x),1-\exp(\min(1,\max(-1,x)))\}$, where $\xi$ acts coordinate-wise on vectors. We obtain the four loss surfaces in Figure 4, which, based on visual inspection, have the same minima. This robustness is a very important property because when the method is applied to a problem without a reference solution, one must be able to trust that the choice of $\psi_{1},\psi_{2},\psi_{3}$ does not need to be tuned based on comparison with a reference solution. Our results indicate a robustness in the choice of $K\geq 2$ and $(\psi_{1},\dots,\psi_{K})$, but it needs to be investigated further, particularly theoretically. The deviation between $Y_{0}$ and the optimized $y_{0}^{*}$ is the discretization error due to the time stepping. This is seen empirically in the convergence plot in Figure 5 for $\xi(x)=x$, showing the experimental convergence order one, i.e., $|Y_{0}-y_{0}^{*}(h)|=\mathcal{O}(h)$. The convergence of $(X,Y,Z)$ as processes after Phase II is shown in Section 4.1.2.
Figure 3: Four MSE curves plotted for $K=1,2,3,4$, with $\psi_{1}=0$, $(\psi_{1},\psi_{2})=(0,b_{1})$, $(\psi_{1},\psi_{2},\psi_{3})=(0,b_{1},-b_{1})$ and $(\psi_{1},\psi_{2},\psi_{3},\psi_{4})=(0,b_{1},-b_{1},-0.5b_{1})$, where $b(x,y,z)=b_{1}(x)+b_{2}(z)$. In Phase II, $y_{0}^{*}$ for $K=1,2,3,4$ are the initial conditions approximated with Phase I. Note that for $K=1$, $y_{0}^{*}=28.44$ as can be visualized in Figure 1. For visualization purposes, we have scaled the curves. For $K=1,2,3,4$, the training times are approximately 400s, 500s, 600s, and 700s, respectively.
Figure 4: Four MSE curves plotted for $\psi_{1}=b_{1}\circ\xi$, $\psi_{2}=-b_{1}\circ\xi$, $\psi_{3}=0$, for $\xi\in\{x,-\sin(x),x\cos(x),1-\exp((x\wedge 1)\vee(-1))\}$, where $\xi$ acts coordinate-wise on vectors. The order coincides with that of the set.
Figure 5: Empirical convergence plot for approximations of $Y_{0}$ using the deep multi-FBSDE method Phase I. Here the case $K=1$ coincides with the original deep FBSDE method.